OpenStudy (anonymous):

∫(5x^3 e^4x )dx evaluate in parts

OpenStudy (anonymous):

5x^3 = u e^4x = dv

OpenStudy (anonymous):

example

OpenStudy (anonymous):

in the picture you have 5x^2 instead of 5x^3 .. so whats the real number?

OpenStudy (anonymous):

thats just an example of how it needs to be worked out from a hw problem

OpenStudy (anonymous):

the problem is 5x^3

OpenStudy (anonymous):

\[∫(5x^3 e^4x )dx \]

OpenStudy (anonymous):

isnt it e^(4x) ?

OpenStudy (anonymous):

yes srry

OpenStudy (anonymous):

\[∫▒(5x^3 e^{4x})x \]

OpenStudy (anonymous):

how did you enter it in wolfram

OpenStudy (anonymous):

With the correct parenthisis.. "integrate (5x^3* e^(4x) )dx"

OpenStudy (anonymous):

how did you get it in parts

OpenStudy (anonymous):

my first post

OpenStudy (anonymous):

for example how would you evaluate this one

OpenStudy (anonymous):

5x^3 = u e^(4x) = dv the formula is: \[\int\limits_{}^{}u*dv = u*v - \int\limits_{}^{}v*du\]

OpenStudy (anonymous):

\[∫_0^4(4e^5+ 3x)dx ]\] jus[t evaluate

OpenStudy (anonymous):

it gives the answer but not the steps

OpenStudy (anonymous):

press the Steps button on Wolfram alpha...

OpenStudy (anonymous):

right it doesnt show that type it and see

OpenStudy (anonymous):

it doesnt let you type integrate for that one and i dont need it in parts just need it to be evaluated

OpenStudy (anonymous):

5x^3 = u e^(4x) = dv the formula is: ∫u∗dv=u∗v−∫v∗du du = 15x^2 v = ∫e^(4x) dx= e^(4 x)/4 now you just need to use this formula untill the "u" disappear.. now your u = 15x^2 dv = e^(4 x)/4 dx derivate the "u" and integrate the "dv" again du = 30x v = ∫ e^(4 x)/4 dx = e^(4 x)/16 apply the formula one last time to eliminate the "u" u = 30x (=) du = 30. now its just a simple integral

OpenStudy (anonymous):

what you mean by evaluating an integral?

OpenStudy (anonymous):

one sec... will post an example

OpenStudy (anonymous):

but you can only evaluate the integral after solving it. First you need to solve it by parts as I already exaplained. when you get the last result 5/128 e^(4 x) (32 x^3-24 x^2+12 x-3) now you can evaluate it..